OpenStudy (anonymous):

The larger leg of a right triangle is 2 feet longer than its smaller leg. The hypotenuse is 4 feet longer than the smaller leg. How many feet long is the larger leg?

OpenStudy (anonymous):

x^2+(x+2)^2=(x+4)^2

OpenStudy (anonymous):

(x-6)(x+2)=0 so x=6 or -2=small leg...it is not -2 cuz its a length...so larger leg=6+2=8

OpenStudy (anonymous):

x=length of small leg

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

yup..the lengths are 6,8 and 10